the area of a rectangle is 675sq.cm. If the width is 1/3 of the length. Find the dimensions of the rectangle????
Answer by nerdybill(7384)

the area of a rectangle is 675sq.cm. If the width is 1/3 of the length. Find the dimensions of the rectangle????
.
Let x = length of rectangle
then from "the width is 1/3 of the length"
x/3 = width of rectangle
.
area = length*width
.
675 = x(x/3)
2025 = x^2
taking the square root of both sides:
45 cm = x (length)
.
width:
x/3 = 45/3 = 15 cm (width)